please help me thanks:
I am two-digit number. the product of my digits is 6. their difference is 1. My tens digit is less than my ones digit. Who am I?

Answer by mananth(16949)

Let the number in the units place be x
then the digit in the tens place will be 6/x

x-(6/x)=1
x^2-6=x
x^2-x-6=0
x^2-3x+2x-6=0
x(x-3)+2(x-3)=0
(x-3)(x+2)=0
x=3 , -2
the other digit =2
So the number is 23